# Regularizing Extrapolation in Causal Inference

### v1

@@ 1 Introduction
A core challenge in observational causal inference and domain adaptation is to adjust data distributions so that features are comparable across distinct groups, such as control and treated groups or source and target populations (Imbens and Rubin, 2015; Farahani et al., 2021). Weighting estimators and linear smoothers, in which the predictions is a weighted average of training outcomes, are widely used for such adjustment; examples include implicit weighting estimators like ordinary least squares (OLS) and random forests and explicit weighting approaches like inverse propensity score weighting (IPW Li et al., 2013) and importance sampling (Thomas and Brunskill, 2017).
@@
An important divide among weighting estimators is whether weights are constrained to be non-negative, such as in traditional IPW, matching (Stuart, 2010), the synthetic control method (Abadie et al., 2010), and stable balancing weights (Zubizarreta, 2015; Ben-Michael et al., 2021a), as well as in the weighting component of popular doubly robust estimators like double machine learning (Chernozhukov et al., 2018). This constraint limits extrapolation and dependence on parametric modeling assumptions, but typically at the cost of worse feature imbalance between re-weighted groups. This imbalance is especially pronounced in high-dimensional settings, when the curse of dimensionality means that positivity is less likely to hold, leading to further bias (D’Amour et al., 2021). By contrast, linear smoothers like OLS and kernel ridge regression allow for arbitrarily negative weights (Robins et al., 2007), which can improve feature imbalance but at the cost of greater model dependence and higher estimator variance. Finally, augmented estimators that combine outcome modeling with explicit weighting strategies can therefore be viewed as performing controlled extrapolation, balancing model dependence against feature imbalance. Pure weighting and pure outcome modeling thus represent two extremes: no extrapolation versus uncontrolled extrapolation.
@@
In this paper, we utilize this geometric perspective to establish a general framework for systematically controlling extrapolation. In particular, we propose a unified approach that directly penalizes the level of extrapolation, replacing the current practice of a hard non-negativity constraint with a soft constraint and corresponding hyperparameter. Unlike prior research on extrapolation in machine learning that emphasizes predictions beyond the observed covariate support, we conceptualize extrapolation through unit weights, a particularly natural framework for handling high-dimensional covariates (Ben-Michael et al., 2021b). Specifically, our contributions are:
@@
  • Bias-bias-variance tradeoff. We propose a framework quantifying a ”bias-bias-variance” tradeoff, decomposing error into bias from distributional imbalance, bias from outcome model misspecification, and estimator variance. This captures key tradeoffs encountered in common causal inference and distribution shift scenarios.
@@ 2.1 Setup and notation
We set up our problem as an instance of prediction/estimation under general distribution shift; the causal inference problems of interest will largely be special cases of these. Consider that we observe $n$ independent and identical draws from source population $P$. For each unit $i\in\{1,\ldots,n\}$, we observe $(X_{i},Y_{i})$, $X_{i}\in\mathcal{X}$ are covariates and $Y\in\mathbb{R}$ is the outcome. We also observe $n_{q}$ iid draws from target population $Q$, with covariates $X\in\mathcal{X}$ but with corresponding missing outcome $Y\in\mathbb{R}$. Finally, define outcome model $\mu(x)=\mathbb{E}[Y\mid X=x]$ and density ratio $dQ/dP(x)$.
For any canonical point in the target population, ${\boldsymbol{x}}^{\star}\in\mathcal{X}$, our goal is to estimate the expected outcome: $\mathbb{E}_{Q}[Y\mid X={\boldsymbol{x}}^{\star}]$. Typically, $\boldsymbol{x}$$^{\star}$ is chosen to be the centroid of the target population (e.g., estimating average treatment effect on treated). The literature typically make following two key assumptions to guide estimation:
@@
    (Conditional Ignorability) $\mathbb{E}_{P}[Y\mid X]=\mathbb{E}_{Q}[Y\mid X]$
@@
Under these restrictions, we can identify the estimand of interest via (1) the outcome function: $\mathbb{E}_{Q}[\mathbb{E}_{P}[Y\mid X]\mid X]$; (2) the weighting function $\mathbb{E}_{Q}\left[\mathbb{E}_{P}\left[dQ/dP(X)Y\right]\mid X\right]$; or (3) via both using the doubly robust formulation. In our setup, we consider situations when A.2. (population overlap) assumption is violated, especially at $x={\boldsymbol{x}}^{\star}$. For regions of $\mathcal{X}$ where A.2. is violated, one needs to rely on parametric assumptions (such as linearity of outcome-covariate relationship) to extrapolate, and identify and estimate the expected outcomes.
@@
Our paper is concerned with parametric model dependence and the bias due to violation of the same. Since we are also focused on linear smoothers, we therefore focus on models that are linear in some features, which are possibly complex functions of the underlying covariates. This is an extremely large model class that ranges from simple linear models to the last layer embedding from a pre-trained large language model. For our setup, we let $\boldsymbol{x}$} be the features in the representation implied by the parametric model, rather than simply the raw covariates.
We further assume:
@@
$\mu$ is Hölder continuous such that $|\mu({\boldsymbol{x}})-\mu({\boldsymbol{x}}^{\prime})|\leq a\cdot\|{\boldsymbol{x}}-{\boldsymbol{x}}^{\prime}\|^{\alpha}$ where, $a>0$ and $\alpha>0$ .
@@ 2.2 Weighting form of causal inference estimators
with weights $w_{\star\leftarrow i}({\boldsymbol{x}}_{i})$, where $\star\leftarrow i$ emphasizes that the weights can depend both on the source covariates ${\boldsymbol{x}}_{i}$ and the target covariates $\boldsymbol{x}$$^{\star}$ (Lin and Han, 2022). When there is no ambiguity, we suppress the dependence on the covariates ${\boldsymbol{x}}_{i}$ and the target $x^{\star}$.
@@
  • Traditional Inverse Propensity Score Weighting. In standard IPW (Rosenbaum, 1987), researchers first estimate a propensity score, $e({\boldsymbol{x}})=\mathbb{P}[1\{i\in P\}\mid{\boldsymbol{X}}_{i}={\boldsymbol{x}}]$ via a binary classifier like logistic regression, and then plug into a known functional form for $dQ/dP(x)$. For example, when estimating the Average Treatment Effect on the Treated, $\hat{w}({\boldsymbol{x}})=\hat{e}({\boldsymbol{x}})/(1-\hat{e}({\boldsymbol{x}}))$. Since $\hat{e}({\boldsymbol{x}})\in(0,1)$, $\hat{w}_{i}({\boldsymbol{x}})>0$ for all $i$.
@@
    $\hat{\mathbf{w}}\in\arg\min_{\mathbf{w}\in\mathcal{W}}\left\|\sum_{i}^{n}w_{i}{\boldsymbol{X}}_{i}-{\boldsymbol{x}}^{\star}\right\|^{2}_{p}+\lambda\|{\boldsymbol{w}}\|^{2}_{2},$ (1)
@@
$\hat{\mu}^{dr}({\boldsymbol{x}}^{\star})$ $=\sum_{i=1}^{N}\hat{w}^{0}_{i}Y_{i}+\left(\hat{m}({\boldsymbol{x}}^{\star})-\sum_{i=1}^{N}w^{0}_{i}\hat{m}({\boldsymbol{x}}_{i})\right)=\hat{m}({\boldsymbol{x}}^{\star})+\sum_{i=1}^{N}\hat{w}^{0}_{i}(Y_{i}-\hat{m}({\boldsymbol{x}}_{i})).$
@@
$\hat{\mu}^{dr}({\boldsymbol{x}}^{*})$ $=\sum_{i=1}^{N}\left(\hat{w}^{0}_{i}+\hat{w}^{\text{adj}}_{i}\right)Y_{i}\qquad\mbox{ where }\;\;\hat{w}^{\text{adj}}_{i}\equiv\hat{\omega}_{i}({\boldsymbol{x}}_{\star})-\sum_{j=1}^{n}\hat{w}^{0}_{j}\hat{\omega}_{i}({\boldsymbol{x}}_{j})$
@@ 3 Regularizing Worst-Case Extrapolation Bias
Second, the difference between $\hat{\mu}({\boldsymbol{x}}^{*})$ and $\hat{\mu}^{\ddagger}({\boldsymbol{x}}^{*})$ is a measure of nonlinearity. In particular, we can decompose $\mu(-{\boldsymbol{x}}_{i})$ as $\delta({\boldsymbol{x}}_{i})-\mu({\boldsymbol{x}}_{i})$ where $\delta({\boldsymbol{x}}_{i})=\left(\mu(-{\boldsymbol{x}}_{i})+\mu({\boldsymbol{x}}_{i})\right)$. In general, $\hat{\mu}({\boldsymbol{x}}^{*})=\hat{\mu}^{\ddagger}({\boldsymbol{x}}^{*})+\sum_{i}|w_{i}|\mathbbm{1}(w_{i}<0)\delta({\boldsymbol{x}}_{i})$; again if $\mu$ is an “odd function”, e.g., $\mu$ is a linear function, then $\delta({\boldsymbol{X}})=0$ because $\mu(-{\boldsymbol{X}})=-\mu({\boldsymbol{X}})$. Thus, $\delta({\boldsymbol{X}})$ is a point-specific measure of nonlinearity in the underlying data generating process.
@@
Although $\delta({\boldsymbol{X}})$ is unknown, we can bound its magnitude using the Hölder continuity assumption: $\|\delta({\boldsymbol{X}})\|\leq 2a\|{\boldsymbol{X}}\|^{\alpha}+2\|\mu({\boldsymbol{X}})\|$. Further, if we assume $\mu(0)=0$, then $\|\delta({\boldsymbol{X}})\|\leq 2a\|{\boldsymbol{X}}\|^{\alpha}$.
The resulting error bound is therefore:
$\left|\mu_{1}({\boldsymbol{x}}^{\star})-\hat{\mu}({\boldsymbol{x}}^{\star})\right|$ $\leq\underbrace{\left|\sum_{i=1}^{n}|w_{i}|\mu({{\boldsymbol{X}}}^{\ddagger}_{i})-\mu({\boldsymbol{x}}^{\star})\right|}_{\text{error in }\hat{\mu}^{\ddagger}({\boldsymbol{x}}^{*})}~+~\underbrace{2a\sum_{i=1}^{n}|w_{i}|\mathbbm{1}(w_{i}<0)\|{\boldsymbol{X}}_{i}\|^{\alpha}}_{\text{error due to nonlinearity}}+\underbrace{\left|\sum_{i=1}^{n}w_{i}\epsilon_{i}\right|}_{\text{noise}}.$ (2)

### v2

@@ 1 Introduction
A core challenge in observational causal inference and domain adaptation is to adjust data distributions so that features are comparable across distinct groups, such as control and treated groups or source and target populations (Imbens and Rubin, 2015; Farahani et al., 2021). Weighting estimators and linear smoothers, in which the prediction is a weighted average of training outcomes, are widely used for such adjustment; examples include implicit weighting estimators like ordinary least squares (OLS) and random forests and explicit weighting approaches like inverse propensity score weighting (Li et al., 2013) and importance sampling (Thomas and Brunskill, 2017).
@@
An important divide among weighting estimators is whether weights are constrained to be non-negative, such as in traditional IPW, matching (Stuart, 2010), the synthetic control method (Abadie et al., 2010), and stable balancing weights (Zubizarreta, 2015; Ben-Michael et al., 2021a), as well as in the weighting component of popular doubly robust estimators like double machine learning (Chernozhukov et al., 2018). This constraint limits extrapolation and dependence on parametric modeling assumptions, but typically at the cost of worse feature imbalance between re-weighted groups. This imbalance is especially pronounced in high-dimensional settings, when the curse of dimensionality means that positivity is less likely to hold, leading to further bias (D’Amour et al., 2021). By contrast, linear smoothers like OLS and kernel ridge regression allow for arbitrarily negative weights (Robins et al., 2007), which can improve feature imbalance but at the cost of greater model dependence and higher estimator variance. Finally, augmented estimators that combine outcome modeling with explicit weighting strategies can therefore be viewed as performing controlled extrapolation, balancing model dependence against feature imbalance. Pure weighting and pure outcome modeling thus represent two extremes: no vs uncontrolled extrapolation.
@@
In this paper, we leverage this geometric perspective to establish a general framework for systematically controlling extrapolation. In particular, we propose a unified approach that directly penalizes the level of extrapolation, replacing the current practice of a hard non-negativity constraint with a soft constraint and corresponding hyperparameter. Unlike prior research on extrapolation in machine learning that emphasizes predictions beyond the observed covariate support, we conceptualize extrapolation through unit weights, a particularly natural framework for handling high-dimensional covariates (Ben-Michael et al., 2021b). Specifically, our contributions are:
@@
  • Bias-bias-variance tradeoff. We propose a framework quantifying a “bias-bias-variance” tradeoff, decomposing error into bias from distributional imbalance, bias from outcome model misspecification, and estimator variance. This captures key tradeoffs encountered in common causal inference and distribution shift scenarios.
@@ 2.1 Setup and notation
To ease exposition, we set up our problem for the causal inference problem of estimating the missing control potential outcome for the Average Treatment Effect on the Treated (ATT). As we note below, however, these results hold for general linear estimands as well as for domain adaptation in ML (Johansson et al., 2022).
For each unit $i\in[n]$, we observe the tuple $(X_{i},Y_{i},Z_{i})$, with covariates $X_{i}\in\mathcal{X}$, outcome $Y_{i}\in\mathbb{R}$, and binary treatment $Z_{i}\in\{0,1\}$. Invoking SUTVA, let $Y_{i}(0)$ and $Y_{i}(1)$ denote the control and treated potential outcomes, respectively, for unit $i$. Our estimand of interest is the ATT, $\mathbb{E}[Y_{i}(1)-Y_{i}(0)\mid Z_{i}=1]$. Since we observe $Y(1)$ for the treated group, the key challenge is to estimate the missing control potential outcome mean, $\mathbb{E}[Y_{i}(0)\mid Z_{i}=1]$. Finally, define the density ratio $dQ/dP(x)$, where $Q$ and $P$ denote the populations of units assigned to treatment and control, respectively. We need the following key assumptions for nonparametric identification:
@@
    (Exchangeability) $\mathbb{E}[Y(0)\mid X,Z=1]=\mathbb{E}[Y(0)\mid X,Z=0]$
@@
In our setup, we consider situations when the population overlap assumption A.2. might be violated. In that case, researchers can instead rely on parametric assumptions on $\mu({\boldsymbol{x}})=\mathbb{E}[Y(0)\mid X_{i}={\boldsymbol{x}}]$, such as linearity, to identify and estimate the expected outcomes. Finally, following Chattopadhyay and Zubizarreta (2023), we will focus on estimating the mean at a target covariate profile, ${\boldsymbol{x}}^{\star}\in\mathcal{X}$, corresponding to our estimand of interest. For the ATT, this profile is simply the mean of the treated population, ${\boldsymbol{x}}^{\star}=\mathbb{E}[X\mid Z=1]$, where $\mu({\boldsymbol{x}}^{\star})=\mathbb{E}[Y(0)\mid Z=1]$.
We set up our problem as an instance of estimating a linear functional. We begin with the fully general setup before specializing to causal inference. Following Bruns-Smith et al. (2023), for each unit $i\in[n]$, we observe the tuple $(X_{i},Y_{i},Z_{i})$, with covariates $X_{i}\in\mathcal{X}$, outcome $Y_{i}\in\mathbb{R}$, and treatment $Z_{i}\in\mathcal{Z}$. The target functional is $\mathbb{E}[h(X_{i},Z_{i},m)]$ for a function $h$, where $m(x,z)=\mathbb{E}[Y_{i}\mid X_{i}=x,Z_{i}=z]$. Many common problems in causal inference and domain adaptation (e.g., Johansson et al., 2022) are special cases of this setup, including counterfactual quantities like the average derivative and the expected policy-specific outcome.
To simplify exposition, we focus on the special case of estimating the missing control potential outcome in the Average Treatment Effect on the Treated (ATT) with binary treatment $Z$. Assuming SUTVA, let $Y_{i}(0)$ and $Y_{i}(1)$ denote the control and treated potential outcomes, respectively, for unit $i$; our estimand is then $\psi=\mathbb{E}[Y_{i}(0)\mid Z_{i}=1]$. Finally, we define the density ratio $dQ/dP(x)$, where $Q$ and $P$ denote the populations of units assigned to treatment and control, respectively. Our goal is to estimate $\psi$ using the observed outcomes of the control units. We need the following key assumptions for nonparametric identification:
  1. A.1.
    (Exchangeability) $\mathbb{E}[Y(0)\mid X,Z=1]=\mathbb{E}[Y(0)\mid X,Z=0]$
  2. A.2.
    (Population overlap) $dQ/dP(x)<\infty$ for all $x\in\mathcal{X}$
In our setup, we consider situations when the population overlap assumption A.2. might be violated. In that case, researchers must rely on parametric assumptions, such as linearity of the outcome-covariate relationship, $\mu({\boldsymbol{x}})=\mathbb{E}[Y(0)\mid X_{i}={\boldsymbol{x}}]$, to extrapolate—and identify and estimate—the expected outcomes. Finally, following Chattopadhyay and Zubizarreta (2023), we will focus on estimating the mean at a target covariate profile, ${\boldsymbol{x}}^{\star}\in\mathcal{X}$, corresponding to our estimand of interest. For the ATT, this profile is simply the mean of the treated population, ${\boldsymbol{x}}^{\star}=\mathbb{E}[X\mid Z=1]$, where $\mu({\boldsymbol{x}}^{\star})=\mathbb{E}[Y(0)\mid Z=1]$.
@@
Since we are focused on linear smoothers, we therefore focus on models that are linear in some features — but which could be possibly complex functions of the underlying covariates. This is an extremely large model class that ranges from simple linear models to the last layer embedding from a pre-trained large language model. For our setup, we let $\boldsymbol{x}$} be the features in the representation implied by the parametric model, rather than simply the raw covariates. We further assume:
@@
$\mu$ is Hölder continuous such that $|\mu({\boldsymbol{x}})-\mu({\boldsymbol{x}}^{\prime})|\leq a\cdot\|{\boldsymbol{x}}-{\boldsymbol{x}}^{\prime}\|^{\alpha}$, with $a>0$ and $\alpha>0$.
@@
General linear estimands.
We can leverage recent work on the Riesz representer (Chernozhukov et al., 2022a) to immediately generalize our results to any linear functional of the data. Following Bruns-Smith et al. (2023), for each unit $i\in[n]$, we observe the tuple $(X_{i},Y_{i},Z_{i})$, with covariates $X_{i}\in\mathcal{X}$, outcome $Y_{i}\in\mathbb{R}$, and treatment $Z_{i}\in\mathcal{Z}$. The target functional is then $\mathbb{E}[h(X_{i},Z_{i},m)]$ for a function $h\in L_{2}$, where $m(x,z)=\mathbb{E}[Y_{i}\mid X_{i}=x,Z_{i}=z]$. Many common problems in causal inference and domain adaptation are special cases of this setup, including counterfactual quantities like the average derivative and the expected policy-specific outcome. Finally, define a feature map $\phi:\mathcal{X}\times\mathcal{Z}\to\mathbb{R}^{d}$; the target feature profile is then $\phi^{*}(x,z)=\mathbb{E}[h(X,Z,\phi)]$. Our results below apply by replacing the simple covariate profile $\boldsymbol{x}$$^{\star}$ with the much more general feature profile $\phi^{\star}(x,z)$.
@@ 2.2 Weighting form of causal inference estimators
with weights $w_{\star\leftarrow i}({\boldsymbol{x}}_{i})$, where $\star\leftarrow i$ emphasizes that the weights can depend both on the source covariates ${\boldsymbol{x}}_{i}$ and the target covariates $\boldsymbol{x}$$^{\star}$ (Lin and Han, 2022). When there is no ambiguity, we suppress the dependence on the covariates ${\boldsymbol{x}}_{i}$ and the target $\boldsymbol{x}$$^{\star}$.
@@
  • Traditional Inverse Propensity Score Weighting. In standard IPW (Rosenbaum, 1987), researchers first estimate a propensity score, $e({\boldsymbol{x}})=\mathbb{P}[Z_{i}=1\mid{\boldsymbol{X}}_{i}={\boldsymbol{x}}]$ via a binary classifier like logistic regression, and then plug into a known functional form for $dQ/dP(x)$. For the ATT, $\hat{w}({\boldsymbol{x}})=\hat{e}({\boldsymbol{x}})/(1-\hat{e}({\boldsymbol{x}}))$; since $\hat{e}({\boldsymbol{x}})\in(0,1)$, $\hat{w}_{i}({\boldsymbol{x}})>0$ for all $i$.
@@
    $\hat{\mathbf{w}}\in\arg\min_{\mathbf{w}\in\mathcal{W}}\left\|\sum_{i=1}^{n}w_{i}{\boldsymbol{X}}_{i}-{\boldsymbol{x}}^{\star}\right\|^{2}_{p}+\lambda\|{\boldsymbol{w}}\|^{2}_{2},$ (1)
@@
$\hat{\mu}^{dr}({\boldsymbol{x}}^{\star})$ $=\sum_{i=1}^{N}\hat{w}^{0}_{i}Y_{i}+\left(\hat{m}({\boldsymbol{x}}^{\star})-\sum_{i=1}^{N}w^{0}_{i}\hat{m}({\boldsymbol{x}}_{i})\right)$
$=\hat{m}({\boldsymbol{x}}^{\star})+\sum_{i=1}^{N}\hat{w}^{0}_{i}(Y_{i}-\hat{m}({\boldsymbol{x}}_{i})).$
@@
$\hat{\mu}^{dr}({\boldsymbol{x}}^{*})$ $=\sum_{i=1}^{N}\left(\hat{w}^{0}_{i}+\hat{w}^{\text{adj}}_{i}\right)Y_{i}\mbox{ where }\;\;\hat{w}^{\text{adj}}_{i}\equiv\hat{\omega}_{i}({\boldsymbol{x}}_{\star})-\sum_{j=1}^{n}\hat{w}^{0}_{j}\hat{\omega}_{i}({\boldsymbol{x}}_{j})$
@@ 3 Regularizing Worst-Case Extrapolation Bias
Second, the difference between $\hat{\mu}({\boldsymbol{x}}^{*})$ and $\hat{\mu}^{\ddagger}({\boldsymbol{x}}^{*})$ is a measure of nonlinearity. We can decompose $\mu(-{\boldsymbol{x}}_{i})$ as $\delta({\boldsymbol{x}}_{i})-\mu({\boldsymbol{x}}_{i})$, where $\delta({\boldsymbol{x}}_{i})=\left(\mu(-{\boldsymbol{x}}_{i})+\mu({\boldsymbol{x}}_{i})\right)$: then $\hat{\mu}({\boldsymbol{x}}^{*})=\hat{\mu}^{\ddagger}({\boldsymbol{x}}^{*})+\sum_{i}|w_{i}|\mathbbm{1}(w_{i}<0)\delta({\boldsymbol{x}}_{i})$. If $\mu$ is an “odd function” (e.g., $\mu$ is linear), then $\delta({\boldsymbol{X}})=0$ because $\mu(-{\boldsymbol{X}})=-\mu({\boldsymbol{X}})$. Thus, $\delta({\boldsymbol{X}})$ is a point-specific measure of nonlinearity in the underlying data generating process.
@@
Although $\delta({\boldsymbol{X}})$ is unknown, we can bound its magnitude using the Hölder continuity assumption: $\|\delta({\boldsymbol{X}})\|\leq 2a\|{\boldsymbol{X}}\|^{\alpha}+2\|\mu({\boldsymbol{X}})\|$.
Further, if we assume $\mu(0)=0$. 111this can also be achieved via centering Then $\|\delta({\boldsymbol{X}})\|\leq 2a\|{\boldsymbol{X}}\|^{\alpha}$.
The resulting error bound is therefore $\left|\mu_{1}({\boldsymbol{x}}^{\star})-\hat{\mu}({\boldsymbol{x}}^{\star})\right|$ less than or equal to
$\underbrace{\left|\sum_{i=1}^{n}|w_{i}|\mu({{\boldsymbol{X}}}^{\ddagger}_{i})-\mu({\boldsymbol{x}}^{\star})\right|}_{\text{error in }\hat{\mu}^{\ddagger}({\boldsymbol{x}}^{*})}+\underbrace{\left|\sum_{i=1}^{n}w_{i}\epsilon_{i}\right|}_{\text{noise}}$ (2)
$+\underbrace{2a\sum_{i=1}^{n}|w_{i}|\mathbbm{1}(w_{i}<0)\|{\boldsymbol{X}}_{i}\|^{\alpha}}_{\text{error due to nonlinearity}}.$